ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

2. Algebra

1991 MS Classification: 18A40, 18B99, 18C15 Linton, Fred E. J., Wesleyan University, Middletown, CT, USA Banach Spaces Monadic over Free Pointed Cones The category of Banach spaces, well known not to be monadic over Sets, turns out to be monadic, after all, over the category of free pointed cones (the Kleisli category for the pointed cone monad). This brief talk will explain why.